OpenStudy (anonymous):

A 1000 kg weather rocket is launched straight up. The rocket motor provides a constant acceleration for 18 s, then the motor stops. The rocket altitude 22 s after launch is 5100 m. You can ignore any effects of air resistance. What was the rocket's acceleration during the first 18 s?

OpenStudy (anonymous):

http://bestdamntutoring.com/Rocket-Propulsion-Kinematics.html This video might help you. You'd just need to switch the variables around.

OpenStudy (anonymous):

d1 = 1/2 *a *(18)^2 Vf = a(18) d2 = Vf*4 -1/2*g*(4)^2 d2 = a(18)*4 -1/2*g*(4)^2 d1 +d2 = 5100 =1/2 *a *(18)^2 + a(18)*4 -1/2*g*(4)^2

OpenStudy (anonymous):

solve for 'a' and stuff

OpenStudy (carson889):

Let d denote the total displacement, d1 denote the distance covered under positive acceleration, d2 denote the distance covered after the motor stops, a1 denote the acceleration in the first 18s. d = 5100m = s1 + s2 d1 = ? d2 = ? a1 = ? a2 = gravity = 9.8 m/s^2 [down] Time under positive acceleration squared = (18s)^2 = 324 Distance traveled under positive acceleration \[d=\frac{ a_{1}*t_{1} ^{2} }{ 2 }=\frac{324a_{1}}{2} =162a_{1}\] Final velocity after acceleration = \[v _{f}= a_{1}*t_{1} + u = 18a_{1} + 0 = 18a_{1}\] \[u = v _{f} = 18a_{1}\] Acceleration under power: \[d _{2}=u*t_{2} + \frac{ a_{2}*t_{2} }{ 2 }=18a_{1}*4 +\frac{a_{2}*4}{2}= 72a_{1} +\frac{-9.8*4}{2}=72a_{1}-19.6 \] \[d=5100 = d _{1}+d_{2}=162a_{1} + 72a_{1}-19.6 = 234a_{1}-19.6\] \[5100+19.6 = 234a_{1} -19.6 + 19\] \[a_{1}=\frac{ 5119.6 }{ 234 } = 21.88\approx21.9\] Therefore, the acceleration during the first 18s was approximately 21.9 m/s^2